the track is a 500m circuit. Ali practices running at a speed of 10kph. Ben practices cycling and travels at 25kph. How long does Ali take to run one lap?
Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)

distance = 500m
speed = 10 kph
(10*1000)/3600 m/s ( 1 km = 1000 m, 1 hour = 3600 sec)
=2.7 m/s
t= d/s
500/2.7
= 185 seconds
Ali takes 185 s for a lap

Answer by greenestamps(13203)



Another wrong answer from tutor @Mananth....!

She converted a fraction to an approximate decimal with only one decimal place, introducing a significant rounding error.

The track is 500m; the speed is 10km/hr = 10000m/hr.

Time is distance divided by speed:

(500m)/(10000m/hr) = 1/20 hr = 3 minutes

ANSWER: 3 minutes for a lap